OpenStudy (anonymous):

Find f(x) if f(2) = 2 and the tangent line at x has slope (x − 1)e^(x^2 − 2x)

OpenStudy (jusaquikie):

calculus 1 or 2?

OpenStudy (anonymous):

f'(x) = (x-1)e^(x^2-2x) Integrate f'(x) to get g(x)= f(x) + k. Evaluate g(2) to obtain the value of k. f(x) = g(x) - k.

OpenStudy (jusaquikie):

yep, lol just like that.